OpenStudy (anonymous):

PLEASE HELP! 3/4 x = 9 4/3 1/4 -3/4 4 Which Number Could be multiplied by each side to produce x = -12

OpenStudy (texaschic101):

multiply by the reciprocal... 3/4x = 9 (4/3)(3/4)x = 9(4/3) 1x (or just x) = 36/3 = 12 check.. 3/4(12) = 9 36/4 = 9 9 = 9 (correct)

OpenStudy (texaschic101):

multiply by 4/3 on both sides

OpenStudy (anonymous):

un hun..

OpenStudy (texaschic101):

oh...wait.....you want x to equal -12....or 12